OpenStudy (lizz123):

Add and subtract polynomials 10(u+v)+8(u-1)-3(u+6) @michele_laino

OpenStudy (lizz123):

What do I have to do first?

oregonduck (oregonduck):

Simplifying -10(u + v) + 8(u + -1) + -3(u + 6) = 0 (u * -10 + v * -10) + 8(u + -1) + -3(u + 6) = 0 (-10u + -10v) + 8(u + -1) + -3(u + 6) = 0 Reorder the terms: -10u + -10v + 8(-1 + u) + -3(u + 6) = 0 -10u + -10v + (-1 * 8 + u * 8) + -3(u + 6) = 0 -10u + -10v + (-8 + 8u) + -3(u + 6) = 0 Reorder the terms: -10u + -10v + -8 + 8u + -3(6 + u) = 0 -10u + -10v + -8 + 8u + (6 * -3 + u * -3) = 0 -10u + -10v + -8 + 8u + (-18 + -3u) = 0 Reorder the terms: -8 + -18 + -10u + 8u + -3u + -10v = 0 Combine like terms: -8 + -18 = -26 -26 + -10u + 8u + -3u + -10v = 0 Combine like terms: -10u + 8u = -2u -26 + -2u + -3u + -10v = 0 Combine like terms: -2u + -3u = -5u -26 + -5u + -10v = 0 Solving -26 + -5u + -10v = 0 Solving for variable 'u'. Move all terms containing u to the left, all other terms to the right. Add '26' to each side of the equation. -26 + -5u + 26 + -10v = 0 + 26 Reorder the terms: -26 + 26 + -5u + -10v = 0 + 26 Combine like terms: -26 + 26 = 0 0 + -5u + -10v = 0 + 26 -5u + -10v = 0 + 26 Combine like terms: 0 + 26 = 26 -5u + -10v = 26 Add '10v' to each side of the equation. -5u + -10v + 10v = 26 + 10v Combine like terms: -10v + 10v = 0 -5u + 0 = 26 + 10v -5u = 26 + 10v Divide each side by '-5'. u = -5.2 + -2v Simplifying u = -5.2 + -2v

OpenStudy (lizz123):

So, for the first part I have to distributive property. Correct?

oregonduck (oregonduck):

yes

OpenStudy (michele_laino):

here you have to apply the distributive propery of multiplication over addition, so we can write: \[\Large \begin{gathered} 10\left( {u + v} \right) + 8\left( {u - 1} \right) - 3\left( {u + 6} \right) = \hfill \\ \hfill \\ = 10u + 10v + 8u - 8 - 3u - 18 = ... \hfill \\ \end{gathered} \] please simplify

OpenStudy (michele_laino):

Hint: next step is: \[\Large \begin{gathered} 10\left( {u + v} \right) + 8\left( {u - 1} \right) - 3\left( {u + 6} \right) = \hfill \\ \hfill \\ = 10u + 10v + 8u - 8 - 3u - 18 = \hfill \\ \hfill \\ = \left( {10 + 8 - 3} \right)u + 10v - 8 - 18 = ... \hfill \\ \end{gathered} \]

OpenStudy (lizz123):

I am so confused

OpenStudy (michele_laino):

why?

OpenStudy (michele_laino):

you have to combine like terms

OpenStudy (lizz123):

 Is that correct?